OpenStudy (anonymous):

A rectangle has perimeter of 40 meters. What length and width yield the maximum area?

OpenStudy (anonymous):

2l + b = 40

OpenStudy (anonymous):

maximize A = lb

OpenStudy (anonymous):

A = l (40-2l)

OpenStudy (anonymous):

l = length, b = width

OpenStudy (anonymous):

only problem is with your reasoning @Kanwar245 is it should be: 2l + 2b = 40 not 2l + b = 40 which means A = l (40-2l) is actually A = l(40-2l)/2 which means A = -l^2 + 40l

OpenStudy (anonymous):

err i meant -l^2 + 20l at the end there. regardless it's easier to solve when it's A = l(20-l)

OpenStudy (anonymous):

thank you!!